OpenStudy (commonlizer):

Two lines, A and B, are represented by equations given below: Line A: y = x − 2 Line B: y = 3x + 4 Which of the following shows the solution to the system of equations and explains why? A (−3, −5), because the point does not lie on any axis B (−3, −5), because the point satisfies both equations C (−4, −8), because the point satisfies one of the equations D (−4, −8), because the point lies between the two axes

OpenStudy (commonlizer):

I already know Line A : y = x - 4 - eqn1 Line B : y = 3x + 4 - eqn2 Plug eqn 1 in eqn 2 x - 4 = 3x + 4 and then x - 4 = 3x + 4 x - 3x = 4 + 4 -2x = 8 x =8/-2 x = -4 and then lug x = -4 in either eqn 1 or eqn 2 y = x - 4 y = -4 - 4 y = -8 I do believe it is C can someone check my answer for me